# Group Projected Subspace Pursuit for Block Sparse Signal Reconstruction: Convergence Analysis and Applications

### v1

@@ Appendix A Uniqueness of block sparse representations
One general characterization involves the compatibility between the kernel space of $\mathbf{A}$ and the block patterns. Indeed, if $\mathbf{A}\widetilde{\mathbf{c}}=\mathbf{A}\bar{\mathbf{c}}$ for some $k$-sparse vectors $\widetilde{\mathbf{c}}\neq\bar{\mathbf{c}}$, then $\widetilde{\mathbf{c}}-\bar{\mathbf{c}}$ is a block $2k$-sparse vector in $\text{ker}(\mathbf{A})$; conversely, any non-zero block $2k$-sparse vector $\mathbf{c}$ admits a decomposition $\mathbf{c}=\widetilde{\mathbf{c}}-\bar{\mathbf{c}}$ where $\widetilde{\mathbf{c}}\neq\bar{\mathbf{c}}$, $|\text{supp}_{\text{G}}(\widetilde{\mathbf{c}})|\leq k$ and $|\text{supp}_{\text{G}}(\bar{\mathbf{c}})|\leq k$, and $\mathbf{c}\in\text{ker}(\mathbf{A})$ implies $\mathbf{A}\widetilde{\mathbf{c}}=\mathbf{A}\bar{\mathbf{c}}$, thus proving the claim. This observation is also made in [15], and we restate it in the following proposition.
@@
We discuss some implications of Proposition 1. First, it indicates that columns within each block need to be linearly independent to ensure uniqueness, i.e., blocks are non-redundant. Second, we deduce that the number of observations $N\geq\sup_{T:|T|\leq 2k}\text{dim}\left(\mathcal{P}_{T}\right)=2kM$ is necessary for unique recovery; this is an instantiation of the minimal sampling requirement for unique recovery from unions of subspaces [31]. Third, the unique block $k$-sparse solution may have multiple $kM$-sparse representations. Indeed, block $k$-sparse vectors are $kM$ sparse vectors but with particular placements of non-zero elements. The kernel of $\mathbf{A}$ being absent of block $2k$-sparse vectors does not mean it does not include $2kM$-sparse vectors. The number of possible distributions of non-zero elements for a $2kM$-sparse vector in $\mathbb{R}^{GM}$ is $\sum_{m=0}^{2kM}{\tbinom{GM}{m}}$, which is much greater than that of a block $k$-sparse vector $\sum_{s=0}^{2k}{\tbinom{G}{s}}(2^{M}-1)^{s}$. This confirms that leveraging the sparsity pattern yields significant dimension reduction.
@@
A detailed proof can be found in [16], where another characterization of uniqueness concerning matrices with almost orthogonal blocks is available. We remark that both Proposition 1 and Proposition 2 provide a macroscopic perspective of the kernel space of $\mathbf{A}$, and the sparsity level within each block is irrelevant.
Leveraging the characterization for unique sparse representations, we show a sufficient condition for unique block sparse representations. Introduced in the classical work by Donoho and Elad [8], the spark of $\mathbf{A}$, denoted by $\text{spark}(\mathbf{A})$, is defined as the smallest integer $m$ such that there exist $m$ columns from $\mathbf{A}$ which are linearly dependent. If $A$ has full column rank, define $\text{spark}(\mathbf{A})=\text{rank}(\mathbf{A})+1$. Employing this concept within each block, we obtain the following result.

### v2

@@ Appendix A Uniqueness of block sparse representations
One general characterization involves the compatibility between the kernel space of $\mathbf{A}$ and the block patterns. Indeed, if $\mathbf{A}\widetilde{\mathbf{c}}=\mathbf{A}\bar{\mathbf{c}}$ for some $k$-sparse vectors $\widetilde{\mathbf{c}}\neq\bar{\mathbf{c}}$, then $\widetilde{\mathbf{c}}-\bar{\mathbf{c}}$ is a block $2k$-sparse vector in $\text{ker}(\mathbf{A})$; conversely, any non-zero block $2k$-sparse vector $\mathbf{c}$ admits a decomposition $\mathbf{c}=\widetilde{\mathbf{c}}-\bar{\mathbf{c}}$ where $\widetilde{\mathbf{c}}\neq\bar{\mathbf{c}}$, $|\text{supp}_{\text{G}}(\widetilde{\mathbf{c}})|\leq k$ and $|\text{supp}_{\text{G}}(\bar{\mathbf{c}})|\leq k$, and $\mathbf{c}\in\text{ker}(\mathbf{A})$ implies $\mathbf{A}\widetilde{\mathbf{c}}=\mathbf{A}\bar{\mathbf{c}}$, thus proving the claim. This observation is also made in [EM09], and we restate it in the following proposition.
@@
We discuss some implications of Proposition 1. First, it indicates that columns within each block need to be linearly independent to ensure uniqueness, i.e., blocks are non-redundant. Second, we deduce that the number of observations $N\geq\sup_{T:|T|\leq 2k}\text{dim}\left(\mathcal{P}_{T}\right)=2kM$ is necessary for unique recovery; this is an instantiation of the minimal sampling requirement for unique recovery from unions of subspaces [LD08]. Third, the unique block $k$-sparse solution may have multiple $kM$-sparse representations. Indeed, block $k$-sparse vectors are $kM$ sparse vectors but with particular placements of non-zero elements. The kernel of $\mathbf{A}$ being absent of block $2k$-sparse vectors does not mean it does not include $2kM$-sparse vectors. The number of possible distributions of non-zero elements for a $2kM$-sparse vector in $\mathbb{R}^{GM}$ is $\sum_{m=0}^{2kM}{\tbinom{GM}{m}}$, which is much greater than that of a block $k$-sparse vector $\sum_{s=0}^{2k}{\tbinom{G}{s}}(2^{M}-1)^{s}$. This confirms that leveraging the sparsity pattern yields significant dimension reduction.
@@
A detailed proof can be found in [EV12], where another characterization of uniqueness concerning matrices with almost orthogonal blocks is available. We remark that both Proposition 1 and Proposition 2 provide a macroscopic perspective of the kernel space of $\mathbf{A}$, and the sparsity level within each block is irrelevant.
Leveraging the characterization for unique sparse representations, we show a sufficient condition for unique block sparse representations. Introduced in the classical work by Donoho and Elad [DE03], the spark of $\mathbf{A}$, denoted by $\text{spark}(\mathbf{A})$, is defined as the smallest integer $m$ such that there exist $m$ columns from $\mathbf{A}$ which are linearly dependent. If $A$ has full column rank, define $\text{spark}(\mathbf{A})=\text{rank}(\mathbf{A})+1$. Employing this concept within each block, we obtain the following result.